OpenStudy (anonymous):

A certain product is packed in cubic boxes that measure 2 feet on a side. If a rectangular shipping container is 6 feet long, 4 feet wide, and 8 feet high, how many of these cubic boxes will be able to fit inside it?

OpenStudy (anonymous):

volume of cubic boxes=2^3=8 feet^3 volume of the rectangular container:6*4*8=192 feet^3 are you able to continue after this?

OpenStudy (anonymous):

YES thank you